# Adaptive Anomaly Recovery for Telemanipulation: A Diffusion Model Approach to Vision-Based Tracking

### v1

@@ I INTRODUCTION
Higher-dimensional data, such as images and videos, provide rich environmental context for complex dexterous manipulation tasks, enabling robotic systems to achieve a more comprehensive understanding of the real-time state space. For instance, in a cube rotation task, raw visual data captures fine-grained finger gaiting movements and intricate object interactions, including changes in contact position and surface sliding. This detailed perception enhances the modeling of manipulation dynamics and supports more reliable control strategies. Consequently, when vision anomalies occur, it is crucial to repair or reconstruct the raw data space to preserve data fidelity. That is to say, for dexterous telemanipulation tasks, it is essential to reconstruct the temporal and continuous video stream during anomalies to maintain operational stability and performance. Traditional video reconstruction techniques, such as optical flow-based tracking and patch-based texture preservation methods  [8, 9], have been widely used to restore missing visual information. However, these approaches often struggle with occlusions, rapid motion, and complex interactions between hands and objects, resulting in inconsistencies and loss of critical details essential for robust manipulation control.
Recent advancements in diffusion-based video generation models have shown notable improvements in frame inpainting, prediction, and reconstruction [10]. These models iteratively refine noisy inputs into high-quality video sequences via large-scale datasets to learn spatiotemporal patterns, enabling smooth and coherent motion. Transformer architectures and attention mechanism further enhance their ability to capture long-range dependencies, ensuring structural consistency and motion continuityn [11]. While these models have proven effective in various visual processing tasks, their potential for predicting the trajectories of manipulated objects—and thereby enhancing the robustness of dexterous telemanipulation—remains largely unexplored.
@@ III-A Anomaly Detection and Video Segmentation
In preliminary experiments, we observed that attempting to use the diffusion model to correct video anomalies with the original video alongside text-based prompts often resulted in unstable reconstructions. This instability is manifested as inconsistencies in object appearance, scale, and timing, particularly in longer video sequences or complex environments. In contrast, providing continuous video—such as frames preceding and following the anomaly—enabled the model to generate more accurate missing frames. Furthermore, the specific frame at which an anomaly is identified significantly influences the reconstruction outcome. For instance, if the input video contains a partial obstruction in motion, the model may inadvertently generate an extended occlusion process rather than restoring the intended scene.
@@
In the FDD approach, the video stream is continuously monitored for anomalies (Algorithm 1). We employ a dual detection strategy that combines frame-difference analysis and ROI tracking. ArUco marker detection enables ROI tracking in this work, but other tracking methods are also applicable. Initially, the system captures the first frame as a baseline and defines an edge-region mask $M$ around the frame borders to detect intrusions (lines 1–3). ArUco marker detection parameters are set (line 4). Each incoming frame undergoes grayscale conversion, Gaussian filtering, and comparison with the baseline (lines 5–9). If pixel differences within $M$ exceed the threshold $P$, the frame is flagged as an intrusion candidate. Concurrently, ArUco markers are tracked (lines 10–11). If they vanish for three or more consecutive frames, an anomaly is confirmed (line 12). Upon detection, the anomalous interval $(S,E)$ is identified (line 15). The video is segmented into $Clip_{A}$ and $Clip_{B}$ to preserve valid frames while discarding corrupted ones (line 16). The processed clips are then sent to the Diffusion model for reconstruction.
@@
13:    if three consecutive frames no ROI detected or $D>P$ then

### v2

@@ I INTRODUCTION
High-dimensional data, such as images and videos, provide rich environmental context for complex dexterous manipulation tasks, enabling robots to understand the real-time state space better. For instance, in a cube rotation task, raw visual data captures fine-grained finger gaiting movements and intricate object interactions, including changes in contact position and surface sliding. This detailed perception enhances the modeling of manipulation dynamics and supports more reliable control strategies. Consequently, when vision anomalies occur, it is crucial to repair or reconstruct the raw data space to preserve data fidelity. Thus, reconstructing the continuous video stream during anomalies is essential for stable telemanipulation. Traditional video reconstruction techniques, such as optical flow-based tracking and patch-based texture preservation methods  [10, 7], have been widely used to restore missing visual information. However, these approaches often struggle with occlusions, rapid motion, and complex interactions between hands and objects, resulting in inconsistencies and loss of critical details essential for robust manipulation control.
Recent advancements in diffusion-based video generation models have shown notable improvements in frame inpainting, prediction, and reconstruction [4]. These models iteratively refine noisy inputs into high-quality video sequences via large-scale datasets to learn spatiotemporal patterns, enabling smooth and coherent motion. Transformer architectures and attention mechanism further enhance their ability to capture long-range dependencies, ensuring structural consistency and motion continuity [13]. While these models have proven effective in various visual processing tasks, their potential for predicting the trajectories of manipulated objects—and thereby enhancing the robustness of dexterous telemanipulation—remains largely unexplored.
@@ III-A Anomaly Detection and Video Segmentation
In preliminary experiments, we observed that using diffusion models with the original video and text prompts often led to unstable anomaly corrections. This instability is manifested as inconsistencies in object appearance, scale, and timing, particularly in longer video sequences or complex environments. In contrast, providing only the continuous video—such as frames preceding and following the anomaly—enabled more accurate reconstructions. Furthermore, the specific frame at which an anomaly is identified significantly influences the reconstruction outcome. For instance, if the input video contains a partial obstruction in motion, the model may inadvertently generate an extended occlusion process rather than restoring the intended scene.
@@
In the FDD approach, the video stream is continuously monitored for anomalies (Algorithm 1). We employ a dual detection strategy that combines frame-difference analysis and ROI tracking. ArUco marker detection enables ROI tracking in this work, though alternative tracking methods are also applicable. Initially, the system captures the first frame as a baseline and defines an edge-region mask $M$ around the frame borders to detect intrusions (lines 1–3). ArUco marker detection parameters are set (line 4). Each incoming frame undergoes grayscale conversion, Gaussian filtering, and comparison with the baseline (lines 5–9). If pixel differences within $M$ exceed the threshold $P$, the frame is flagged as an intrusion candidate. Concurrently, ArUco markers are tracked (lines 10). If they vanish for $k$ or more consecutive frames, an anomaly is confirmed (line 11–12). Upon detection, the anomalous interval $(S,E)$ is identified (line 15). The video is segmented into $Clip_{A}$ and $Clip_{B}$ to preserve valid frames while discarding corrupted ones (line 16). The processed clips are then sent to the Diffusion model for reconstruction. The parameter values used in our implementation are $B=50$, $T=100$, $P=1000$, and $k=3$.
@@
13:  if no ROI is detected for $k$ consecutive frames or $D>P$ then